OpenStudy (anonymous):

A 68-mile trip uses two gallons of gas. How many gallons would you use on a 110-mile trip? Round to the nearest tenth of a gallon.

OpenStudy (anonymous):

7480 gallons

OpenStudy (anonymous):

Ithink your a little off

OpenStudy (anonymous):

nopee calculator is right

OpenStudy (anonymous):

hld on lol

OpenStudy (anonymous):

3.6 i thinkk

OpenStudy (anonymous):

srry i was doing a diff problem lol

OpenStudy (anonymous):

where you from?

OpenStudy (anonymous):

who?

OpenStudy (anonymous):

florida

OpenStudy (anonymous):

oh nvm you looked like a different chelsea

OpenStudy (anonymous):

whyy u ask

OpenStudy (anonymous):

cause you looked like someone else i knew

OpenStudy (anonymous):

\[\frac{110}{\frac{68}{2}}\text{=}\frac{55}{17}=3.24 \text{ gallons}\]

OpenStudy (anonymous):

okk